OpenStudy (anonymous):

if I have a graph of several exponential functions, what will a logarithmic scale of the y-axis show me?

OpenStudy (jamesj):

Well, if y = C exp(ax), i.e., y an exponential function. Then if now take the log of the y-axis, we have \[\ln y = \ln C + a x\] And that is the equation of ... what?

OpenStudy (jamesj):

So be absolutely clear, write y' as the new y axis (not the derivative). Then y' = ax + c, where c = ln C. What does the graph of that look like?

OpenStudy (anonymous):

this will look like a linear function, but I'm actually more interested in what a logarithmic scale will tell me about the relation of the exponential functions

OpenStudy (anonymous):

if the slope of one function (in log scale) is smaller than another, will this mean that the exponent of that function is smaller than the one with a higher slope?

OpenStudy (jamesj):

The slope of the lines is now the index in the exponential, and the y'-axis intercept is the constant. So if you have two lines that are parallel, they must have the same index in the exponential.